OpenStudy (anonymous):

work out 8d-4d-6d-8=2d

OpenStudy (anonymous):

8d-4d-6d-8 = 2d <----- Adding like terms... -2d-8 = 2d -2d -2d -------------- -4d-8 = 0 +8 +8 ------------ -4d = 8 /-4 /-4 ------------ d = -2

Directrix (directrix):

@chandhooguy OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Ok @Directrix I am Sorry